Munn Inc. reported other noncurrent asset account balances on December 31, 2020, as follows.

Patent $384,000
Accumulated amortization (48,000)
Net patent $336,000

Transactions during 2021 and other information relating to Munn’s other noncurrent assets include the following.

1. The patent was purchased from Grey Company on January 2, 2019, when the remaining legal life was 16 years. On January 1, 2021, Munn determined that the remaining useful life of the patent was only eight years from the date of its acquisition.

2. On January 1, 2021, in connection with the purchase of a trademark from Cody Corp., the parties entered into a noncompetition agreement. Munn paid Cody $1,600,000, of which 75% related to the trademark and 25% reflected Cody’s agreement not to compete for a period of five years in the line of business covered by the trademark. Munn considers the life of the trademark to be indefinite.

3. On January 1, 2021, Munn acquired all the noncash assets and assumed all liabilities of Amboy Company at a cash purchase price of $2,400,000. Munn determined that the fair value of the identifiable net assets acquired in the transaction is $1,600,000.

4. Munn incurred the following research and development costs in 2021.

Salaries and related expenses related to development of patents $250,000
Supplies 30,000
Allocated facility costs 40,000

Required

a. Prepare a schedule of amortization for 2021 relating to Munn’s intangible assets assuming straight-line amortization.

2021 Amortization Expense
Patent
Noncompete agreement
Total

b. Prepare the balance sheet presentation of intangible assets for Munn on December 31, 2021, and the income statement presentation for the year ended December 31, 2021.

Munn Inc.
Balance Sheet
December 31, 2021
Intangible assets
Patent, net
Trademark
Noncompete agreement, net
Goodwill
Munn Inc.
Income Statement
For the year ended December 31, 2021
Other expenses
Amortization expense
Research and development expense
Answer #1

a.

2021 Amortization Expense
Patent 56000
Noncompete agreement 80000
Total 136000

The book value of the patent on January 1, 2021, $336000 will be amortized over its remaining useful life which is 8 - 2 = 6 years.

Amortization expense for patent = $336000/6 = $56000

Amortization expense for noncompete agreement = (25% x $1600000)/5 = $400000/5 = $80000

b.

Munn Inc.
Balance Sheet
December 31, 2021
Intangible assets
Patent, net 280000
Trademark 1200000
Noncompete agreement, net 320000
Goodwill 800000

Patent, net = $336000 - $56000 = $280000

Trademark = 75% x $1600000 = $1200000

Noncompete agreement, net = $400000 - $80000 = $320000

Goodwill = $2400000 - $1600000 = $800000

Munn Inc.
Income Statement
For the year ended December 31, 2021
Other expenses
Amortization expense 136000
Research and development expense 320000

Research and development expense = $250000 + $30000 + $40000 = $320000